OpenStudy (anonymous):

Find the slope of the line that contains these points. (3, 2) (5, 12)

OpenStudy (anonymous):

does anyone know the answer

terenzreignz (terenzreignz):

Plenty of people here know the answer... and if you 'listen' well, you will too :D If you're given two points, (a,b) and (c,d), the slope of the line through them is given by this equation... \[\Large m = \frac{d-b}{c-a}\]

terenzreignz (terenzreignz):

So, in this case, your (a,b) is (3,2) and your (c,d) is (5,12) Just plug in and solve for the slope (m).

terenzreignz (terenzreignz):

Confused?

OpenStudy (anonymous):

no thanks I got it

terenzreignz (terenzreignz):

Okay... what's your answer?

OpenStudy (anonymous):

1,-7

terenzreignz (terenzreignz):

Oh... no, those aren't the differences you need... When you have an ordered pair, (x,y) we call the left coordinate the x-coordinate and the right coordinate the y-coordinates, right? Your task is not to subtract the x and y coordinates, but rather, Get the difference of their y coordinates, and get the difference of their x coordinates... can you do that?

OpenStudy (anonymous):

so would that make it their oppisites

terenzreignz (terenzreignz):

I can see how that might be confusing... let me put up an example... Say we need the slope of the line through the points (1 , 7) and (3 , 21) Their x-coordinates are 1 and 3 their y-coordinates are 7 and 21. Get the difference of their x-coordinates, 1-3 = -2 AND the difference of their y-coordinates, 7-21 = -14 The slope is "difference of y-coordinates divided by the difference of x-coordinates" So, it's -14/-2 = 7

OpenStudy (anonymous):

thanks for mapping it out for me it helped